A drug test is accurate 97% of the time. If the test is given to 2300 people who have not taken drugs, what is the probability that at least 71 will test positive?
Answer #1

By using normal approximation:

Mean = np = 2300*0.03 = 69

Standard deviation = npq = 2300 * 0.97 * 0.03 = 8.1811

Hence,

P(Atleast 71)

= P(X > 70.5) [Continuity Correction]

= P(z > (70.5 - 69)/8.1811)

= P(z > 0.18)

= 0.4273